Alex drove 500 miles in 15 days.
At the same rate, how long would it take Alex to drive 5,000 miles?

Answer by checkley75(3666)

15/500=X/5,000
500X=15*5,000
500X=75,000
X=75,000/500
X=150 DAYS TO DRIVE 5,000 MILES.